diffrentitate this equation (x2 – 7)2........   please explan it’s urgent
                                              4x

Grade:12

1 Answers

Arun
5 years ago
 
Dear Shubham
 
if we differentiate form of f/g 
then,
 
d(f/g) / dx = (f * g’ – g * f’ )/ g^2
 
Apply this
 
dy/dx = (4 (x^2 – 7)^2 – 4x * 2 (x^2 – 7) * 2x )/ 16 x^2
 
Hope it helps
 
Regards
Arun (askIITians forum expert)